Suppose that a foreign company withdraws money from its account in a foreign country, buys 1 million US dollars and deposits this into the US banking system. The target reserve ratio is 10% and there is no currency drain from the banking system.
ALL BANKS HAVE
Reserves $90 million,
deposits: 900 million
Loans: 710 million
securities 100 million

What are the target reserves and what are the excess reserves?
After 1 million deposit what are the target and excess reserves?
All excess reserves are used to expand loans, the final increase in the money supply will be _?_ times the new 1 million deposit, which is equal to _?_
What is the final increase in loans?
1
Expert's answer
2018-12-14T14:10:10-0500

The target reserves are 900*0.1 = $90 million.

The excess reserves are zero, because the target reserves equal required reserves.

After 1 million deposit the target reserves are 901*0.1 = $90.1 million and excess reserves are zero again.

If all excess reserves are used to expand loans, the final increase in the money supply will be 1/0.1 = 10 times the new 1 million deposit, which is equal to $10 million.

The final increase in loans is $10 million.